OpenStudy (anonymous):

Can someone walk me through the indefinite integral of: t(ln(t+1))?

OpenStudy (anonymous):

Rewritten: \[\int\limits_{ }^{} t \ln(t+1) dt\]

OpenStudy (turingtest):

Looks like you could use integration by parts

OpenStudy (anonymous):

Yes that's what we are supposed to be exercising, but I don't understand how to do this. What would I use for u, and what would I use for dv?

OpenStudy (turingtest):

\[\int t\ln(t+1)dt\]\[u=\ln(t+1)\implies du=\frac{dt}{t+1}\]\[dv=tdt\implies v=\frac12t^2\]

OpenStudy (turingtest):

nah maybe not, let me try on paper...

OpenStudy (turingtest):

oh yes, it will work the way I did it :)

OpenStudy (turingtest):

presumably you got to\[\int t\ln(t+1)dt=\frac t{t+1}-\frac12\int\frac{t^2}{t+1}dt\]and got stuck?

OpenStudy (anonymous):

Haha I actually wasn't sure what to pick for u or uv, and why.

OpenStudy (anonymous):

dv* not uv

OpenStudy (turingtest):

well you have two terms, \(t\) and \(\ln(t+1)\) so one must be u and one must be dv...

OpenStudy (turingtest):

you should think to yourself "I don't know how to integrate \(\ln(t+1)\), so it doesn't make a lot of sense to use it for dv since I will have to integrate that"

OpenStudy (turingtest):

that means it must be u, which means t must be dv reasonable?

OpenStudy (anonymous):

Okay and ln(t+1) is very easy to find the derivative of. So I should pick it as u

OpenStudy (turingtest):

you actually could integrate it, and I bet this would work just fine that way though ;) bvut I think it is easier to reason it my way yes u=ln(t+1)

OpenStudy (turingtest):

but*

OpenStudy (anonymous):

ok so it would look like this:

OpenStudy (anonymous):

(t^2 / 2)(ln(t+1) - integral(t/(t+1)) ?

OpenStudy (anonymous):

sorry if that looks messy

OpenStudy (turingtest):

oh we both messed up :/

OpenStudy (turingtest):

\[\int t\ln(t+1)dt=\frac12t^2\ln(t+1)-\frac12\int\frac{t^2}{t+1}dt\]

OpenStudy (anonymous):

Where did the 1/2 come from infront of the integral

OpenStudy (turingtest):

\[u=\ln(t+1)\implies du=\frac{dt}{t+1}\]\[dv=tdt\implies v=\frac12t^2\]\[\int udv=uv-\int vdu\]\[\int t\ln(t+1)dt=\frac12t^2\ln(t+1)-\frac12\int\frac{t^2}{t+1}dt\]

OpenStudy (turingtest):

it came from the fact that v=t^2/2

OpenStudy (anonymous):